Circular No. 9037

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 G1 (STEREO) K. Kadota (Ageo, Japan) has reported astrometry on Apr. 9.8 UT

with a 25-cm reflector, which shows that the prediction on IAUC 9036 was off by ∼ 4′. He found the comet to have total mag 10.6, a coma diameter of 4′.5, and strong central condensation; his astrometry and the following revised orbital elements appear on MPEC 2009-G32:

T = 2009 Apr. 15.093 TT  ω = 173.^{o}823
Ω = 120.657 }2000.0
q = 1.12993 AU  i = 108.091

COMET C/2009 F6 (YI-SWAN) S. Korotskiy, Moscow, reports that he found a prediscovery image of

this comet (cf. IAUCs 9034, 9035) on an exposure (limiting mag ∼ 13) taken with a Canon EOS 20D digital camera (+ 50-mm f/4 lens) at the Kazan State University observatory at Karachay-Cherkessia, as given below:

2009 UT  α_{2000}  δ_{2000}  Mag.
Mar. 25.06024  22^{h}28^{m}28.^{s}8  +48^{o}45′14′′  10.7

COMETS C/2003 Q1, Q6; C/2008 X6; C/2008 Y10, Y11 (SOHO) Further to IAUC 9036, additional near-sun presumed comets have been

found on SOHO website images (with the initial observations tabulated below). C/2008 Y10, a Kreutz sungrazer, was stellar in appearance (mag ∼ 6) in C3 images; in C2 images, it was teardrop-shaped and slightly diffuse. C/2008 Y11, a member of the Marsden group, was tiny and stellar in appearance (mag ∼ 7). R. Kracht proposed that C/2008 Y11 may be identical with C/2003 Q6 (cf. IAUC 8339), and C/2008 X6 (cf. IAUC 9030) with C/2003 Q1 (cf. IAUC 8339); linkages by B. G. Marsden (cf. MPEC 2009-F81) suggest that these two comets (both with P ≃ 5.3 yr) may have separated from each other around their last return to perihelion (1998 Apr. 26–29), though Kracht was unable to find any trace of them in C2 data from 1998 Apr. 25.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 Y10 Dec. 21.988 18^{h}09.^{m}0 −27^{o}29′ C3/2 JR 2009-F17
C/2008 Y11  22.288 17 59.0 −22 54 C2  RK 2009-F17
2009 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green